OpenStudy (anonymous):

Through (-4, -5) parallel to y=3/5x-1 with steps?

OpenStudy (anonymous):

y+5 = 3(x+4)/5

OpenStudy (anonymous):

let equation of line be y=mx+c m is slope of line and c is some constant since line is parallel to y=0.6x-1 so, slope of y=mx+c is also 0.6 i.e. m=0.6 line passes through (-4,-5) this point must satisfy our line hence we can find c = -2.6 y=0.6x-2.6 (him1618's answer)